The monthly demand for ice cans being manufactured by Mr. Alarde is 3200 pieces.  With a manual operated guillotine, the unit cuffing cost is P25.00. An electrically operated  hydraulic guillotine was offered to Mr. Alarde at a price of P275,000.00 and which cuts by  30% the unit cuffing cost. Disregarding the cost of money, how many months will Mr. Alarde be able to recover the cost of the machine if he decides to buy now? 



1
Expert's answer
2021-12-07T18:52:14-0500

Monthly saving = 3,200 x P25 x 0.3 = P24,000

Payback period (PBP) measures the time by when cost is recovered by savings, and is the month by when cumulative net cash flow is zero.



PBP lies between months 11 and 12.

PBP (Months) = 11 + (Absolute value of cumulative cash flow, month 11 / Cash flow, month 12)

= 11 + (11,000 / 24,000) = 11 + 0.46

= 11.46